Grade 12th passPhysical Chemistry

What is mass percent of oxygen in Al2 (SO4)3.18H2O?(M=666.43 g/mol)?

Anjali
9 Years agoGrade 12th pass

1 Answer

Nanda Vinod
9 Years ago
mass percent of an element in a compound= mass of the element present ina given mass of the compound*100                                                                                                      mass of the compound taken
1 mole of the given compund weighs 666g and i contains 12+18=30 moles of O which weighs 30*16=480g
mass of O=480g/mol.
mass % of O =480*100/666=72.1%